help me with question 4 all pleasee..please show your step :')with three goods. 1 .*y, and and 3. There are 200 4. (20 points) Consider a pure exchange economy with three goods, consumers
Answer #1

1. Let p = 1, & P2, P3 → Prices of Good 2 & 3 endowment Income MA = 1P,+ 1 P2 + opz = 1+P2 MB = opit IP2 + 1P3 = P2+83 I UCX1then [X1B = $s (Peats) ja Ves = (Pzf9 | .] *3* = + (fotfa) ] (1) Now at walsasian eam excess dem and in all Mkts are cleared